Several studies have shown that women with many children are less likely to get ovarian cancer. In a new study, data are collected from 25 women ages 40-49 with ovarian cancer. The mean parity (number of children) of these women is 1.8 with standard deviation 1.2. Suppose the mean number of children among women in the general population in this age group is 2 .5. ovarian cancer have fewer children e u than women in the general population in the same age group? 2. Perform the toat using the critical-value method. 3. What is the p-value based on the test? 4. What do you conclude from this study?
Answer #1

Given Information,

n=25, , , , ar{x} = 1.8, , , , s= 1.2

Our hypothesis is,

2.5

H1 : μ < 2.5 (Left tailed test)

1.  Since, sample size is small (n<30) and population standard deviation (sigma) is unknown, so t-test is use here.

2. Test statistics:

18-25-11-29166671-2.9167 sMn = 1.2 25 0.24

Test statistics is 2.9167

The level of significance (alpha) is not given in the problem, so if a-0.05

Degree of freedom = n-1 =25-1= 24

Therefore, critical t value at 0.05 level of significance for 24 degree of freedom is 1.711 (From t- table)

Since, calculated t is greater than the critical t value, i.e.,

tcal 1.711, Ho is rejected .

Therefore, we conclude that there is enough evidence to support the claim that the women with ovarian cancer have fewer children than women in the general population in the same age group.

3. P- value: P value is calculated in R as:

(1-pt(2.9167,24)) = 0.00378

4. Conclusion: Since, p value is less than the specified level of significance, i.e.,

  < 0.05, Ho is rejected

Therefore, we conclude that there is enough evidence to support the claim that the women with ovarian cancer have fewer children than women in the general population in the same age group.